OpenStudy (anonymous):

At a laundromat, there are 12 washing machines available for the people to use. On a particular day, a total of 42 people came and did their laundry where each person uses exactly one washing machine. Show that if no washing machine is used by more than 6 people, then at least 5 washing machines are used by at least 3 people.

OpenStudy (rational):

pigeonhole principle ?

OpenStudy (anonymous):

yes but i confused. there is a hint : Consider when the maximum number of people uses 4 washing machines. Show that that there is one more washing machine being used by at least 3 people.

OpenStudy (anonymous):

pigeonhole = machine pigeon = people right ??

OpenStudy (rational):

Yes. say the first four washing machines are used by maximum number of people : 4*6 = 24 (24 people used the first 4 washing machines.) (24 pigeons in first 4 holes)

OpenStudy (rational):

how many more washing machines (holes) are there ? how many more people(pigeons) are there ?

OpenStudy (anonymous):

8 washing machines and 18 people. 15 people use 5 washing machine, then 3 people left. thats why 1 more machine use by at least 3 people. yes or no ?

OpenStudy (rational):

18 pigeons and 8 holes by pigeonhole principle, one hole must contain at least 3 pigeons

OpenStudy (rational):

18/8 = 2.25 so there will be few pigeons left after assigning 2 pigeons to each hole

OpenStudy (rational):

that means "one" of the 18 washing machines is used by at least 3 ppl

OpenStudy (anonymous):

okay.. thats mean each of 7 holes used by 2 pigeon , so 14 pigeons. 18-14 =4 pigeons left. 4 = at least 3 in 1 more pigeonhole.. right ?

OpenStudy (rational):

thats right, but if you look at the division : 18/8 clearly thats more than 2, that means one of the hole must contain more than 2 pigeons

OpenStudy (anonymous):

yes.. okay..thank you very much for your help.. :)

OpenStudy (rational):

np :) do you get how to conclude "at least 5 washing machines are used by at least 3 people."

OpenStudy (anonymous):

i dont know .. can you explain it ?

OpenStudy (rational):

we just showed that at least one of the last 8 washing machines is used by at least 3 ppl right ?

OpenStudy (anonymous):

im not sure. but if you say so..then its right. let me try with my friends first. if not okay i'll ask you later. :)

OpenStudy (rational):

ok sure take your time :)

OpenStudy (rational):

``` 18 pigeons and 8 holes by pigeonhole principle, one hole must contain at least 3 pigeons ``` this shows that one of the 18 washing machines must be used by at least 3 ppl.

OpenStudy (anonymous):

okay.. i'll try it.. thanks.. :)

OpenStudy (rational):

yw:)